OpenStudy (anonymous):

help i put B am i right? What is the y-intercept of the line 3x + 4y = 16? (Points : 2) 16 4 3

OpenStudy (anonymous):

@calculusfunctions

OpenStudy (anonymous):

If 4 is b, then you are right

OpenStudy (anonymous):

thnks :)

OpenStudy (calculusfunctions):

YES B is correct!

OpenStudy (anonymous):

wat about this one i put A . What is the slope of the line 2x – 5y = 12 ? (Points : 2) 2

OpenStudy (anonymous):

A=2/5 B=5/2 C=2 D=-12/5

OpenStudy (calculusfunctions):

NO! If you have Ax + By = C, then the slope is given by -A/B.

OpenStudy (calculusfunctions):

Then you are correct, the answer is A.

OpenStudy (anonymous):

so its D>?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

is this true or false i put true This figure shows the correct way of constructing a line perpendicular through a given point.

OpenStudy (calculusfunctions):

When we're both type at the same time, misunderstandings are likely.

OpenStudy (calculusfunctions):

What figure?

OpenStudy (calculusfunctions):

OH!

OpenStudy (anonymous):

i put true am i right?

OpenStudy (anonymous):

@calculusfunctions

OpenStudy (anonymous):

???

OpenStudy (calculusfunctions):

Do you mean a line perpendicular to a line through a given point? I was taking my time because a line perpendicular to a point makes no sense.

OpenStudy (calculusfunctions):

A point has no dimensions so how can anything be perpendicular to it? Please always double check that you have written the questions correctly.